OpenStudy (anonymous):

Find the product of (x + 7)2. x2 - 14x + 49 x2 + 14x + 49 x2 - 49 x2 + 49

OpenStudy (anonymous):

@girlover @bpstation @wyatthelean96 @ajprincess @johnweldon1993 @undeadknight26 @prowrestler @gibran97

OpenStudy (girlover):

B

OpenStudy (haseeb96):

@girlover first tell him the formula then solve his question then tell him answer

OpenStudy (anonymous):

Find the product of (3x + 7y)(3x - 7y). 9x2 - 42xy + 49y2 9x2 + 42xy + 49y2 9x2 - 49y2 9x2 + 49y2

OpenStudy (anonymous):

how about this one?

OpenStudy (haseeb96):

(a+b)^2 = a^2 + 2ab +b^2 so in the question (x+7)^2 = x^2 + 2 (x) (7) + 7^2 =x^2 +14x +49 is the correct answer

OpenStudy (haseeb96):

(a+b)(a-b) = (a^2 - b^2 ) (3x + 7y)(3x -7y) = ( 9x^2 - 49y^2)

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

Find the product of (x - 7)2. x2 - 14x + 49 x2 + 14x + 49 x2 - 49 x2 + 49

OpenStudy (haseeb96):

(a-b)^2= a^2 -2ab +b^2 use this formula dn get the answer yur self and tell me the answer

OpenStudy (anonymous):

D?

OpenStudy (haseeb96):

A is the correct answer if u like my explaination then click on best response

OpenStudy (anonymous):

no ty and ty